OpenStudy (anonymous):

Please help! I need to finish this by tonight. Find the amplitude, period, phase shift, and vertical translation for these two equations: 1. y= -sin(x - pi/4) + 2 2. y= 2cos(2(pi)(x))

OpenStudy (whpalmer4):

Amplitude is half the difference between maximum and minimum values. For example, the amplitude of \(y=\sin x\) is \(\frac{1}{2}(1-(-1)) = 1\). Period is the smallest interval over which the function repeats. For \(y=\sin x\) the period is \(2\pi\). If you have a function \(y = \sin Bx\) the period will be \(\frac{2\pi}B\). Phase shift is the value added to the argument of the periodic function. \(y=\sin (2x+3\pi)\) has a phase shift of \(3\pi\). Vertical translation is any offset added to the result of the periodic function. For example, \(y=\sin x + 1\) has a vertical translation of 1, and oscillates between 0 and 2 instead of -1 and 1.

OpenStudy (whpalmer4):

Some examples: First one shows the effects of changing amplitude, phase shift, and translation. Second one shows changing period, and then a combination of amplitude, phase shift and vertical translation.

OpenStudy (anonymous):

Shouldn't the one labeled as \[y=2\sin(2x+\frac{ \pi }{ 3 })\] in sines2.png be \[y=2\sin(2x+\frac{ \pi }{ 3 })+1?\]

OpenStudy (whpalmer4):

Nuts, I must have brushed against the delete key after typing it...you're right!

OpenStudy (anonymous):

No problem! How did you make those graphs so fast btw?

OpenStudy (whpalmer4):

Mathematica is wonderful, so long as you can figure out how to tell it what you want to do!

OpenStudy (anonymous):

ooh, i think i got it. 1. amplitude:1 period:2pi phase shift: -pi/4 vertical translation: 2

OpenStudy (whpalmer4):

Yes, I believe those are correct.

OpenStudy (anonymous):

Alrighty, thanks so much!

OpenStudy (whpalmer4):

How about the second problem?

OpenStudy (anonymous):

Oh goodness, how could I forget about that?

OpenStudy (anonymous):

Hold on a second.

OpenStudy (whpalmer4):

drum roll please.... :-)

OpenStudy (anonymous):

2. amplitude: 2 period: 1

OpenStudy (whpalmer4):

phase shift and vertical translation?

OpenStudy (anonymous):

but for phase shift and vertical translation, I'm not so sure.

OpenStudy (whpalmer4):

how about none? does none strike your fancy?

OpenStudy (anonymous):

hmm. none. can you elaborate on that?

OpenStudy (whpalmer4):

There is no phase shift, because you don't have \(2\pi x + \phi\) as the argument to \(\cos\). There's no vertical translation, because you don't have anything added on to the result of \(2\cos (2\pi x)\) either.

OpenStudy (whpalmer4):

If you had all the options, the formula would look like \[y = A \sin (Bx + \phi) + C\] where \(A\) controls amplitude, \(B\) controls period (or frequency), \(\phi\) is the phase shift, and \(C\) is vertical translation or offset.

OpenStudy (anonymous):

Ahh, that's why I was so utterly stuck in that quagmire. I was overthinking!

OpenStudy (anonymous):

Thank you so much!~

OpenStudy (whpalmer4):

It's not quite as helpful as it might be, because it doesn't show you the function you've chosen, but check out http://demonstrations.wolfram.com/SineAndCosineGraphGenerator/ (I think you have to install the Wolfram CDF player, there should be a link on the page) You can play with vertical stretch, phase shift, vertical shift, period with slides and observe the result.

OpenStudy (anonymous):

Okie doke! I'll check it out.

OpenStudy (whpalmer4):

Also very interesting if you're curious about possible applications for the magic sin and cos functions appropriately squashed, stretched, shifted, etc. is http://demonstrations.wolfram.com/FourierSeriesOfSimpleFunctions/ Fourier Series allow you to build up just about any waveform by combining the right selection of sines and cosines. This turns out to be an extremely powerful concept in many fields...